The average age of 10 children is 9 years 9 months. The average age of the first 9 children is 8 years 11months.what is the age of the 10th child?
Answer by Reken R Maikama(2)

x/10=9.75 then x=390/4,now ave of 9 students y/9=8.9 then y=963/12. age of the 10th student will be x-y=17.25